While finding the amount of seed needed to plant his three square wheat fields, Hank observed that the side of one field was 1 kilometer longer than the side of the smallest field and that the side of the largest field was 3 kilometers longer than the side of the smallest field. If the total area of the three fields is 38 square kilometers, then what is the area of each field?

Answer by checkley77(12844)

Let x be the side of the smaller fild.
x+1 is the side of the larger field.
x+3 is the side of the largest field.
x^2+(x+1)^2+(x+3)^2=38
x^2+x^2+2x+1+x^2+6x+9=38
3x^2+8x+10-38=0
3x^2+8x-28=0
(3x+14)(x-2)=0
x-2=0
x=2 km. is the side of the smallest square.
2+1=3 km for the side of larger square.
2+3=5 km is the side of the largest square.
Proof:
2^2+362+5^2=38
4+9+25=38
38=38